# Residual lifetime prediction for heterogeneous degradation data by Bayesian semi-parametric method

### v1

@@ Abstract
Degradation data are considered for assessing reliability in highly reliable systems. The usual assumption is that degradation units come from a homogeneous population. But in presence of high variability in the manufacturing process, this assumption is not true in general; that is different sub-populations are involved in the study. Predicting residual lifetime of a functioning unit is a major challenge in the degradation modeling especially in heterogeneous environment. To account for heterogeneous degradation data, we have proposed a Bayesian semi-parametric approach to relax the conventional modeling assumptions. We model the degradation path using Dirichlet process mixture of normal distributions. Based on the samples obtained from posterior distribution of model parameters we obtain residual lifetime distribution for individual unit. Transformation based MCMC technique is used for simulating values from the derived residual lifetime distribution for prediction of residual lifetime. A simulation study is undertaken to check performance of the proposed semi-parametric model compared with parametric model. Fatigue Crack Size data is analyzed to illustrate the proposed methodology.
@@ 1 Introduction
For some highly reliable engineering products, it is difficult to estimate reliability due to the fact that products take too long time to fail. In this context, degradation measures taken over time are used to estimate reliability. The evolution of the degradation measures may be observed using sensor technology through a procedure known as Condition Monitoring (See Nelson [1]). Some examples of degradation data include vibration signals for monitoring excessive wear induced in rotating machinery, acoustic emissions for monitoring crack propagation, temperature changes and oil debris for engine lubrication, decrease of brightness of light bulbs, etc. Inferences on lifetime distribution and residual lifetime of a product can be done by modeling the underlying degradation mechanism that represents the evolution of degradation resulting failure. There have been a number of works on degradation modeling, see for example Lu and Meeker [2], Padgett and Tomlinson [3], Muller and Zhang [4] and Park and Padgett [5]. Degradation data provide more information than the bare lifetimes and hence precise inferences can be made as discussed in Lu and Meeker [2]. The modeling of degradation data can be considered by different processes. The commonly used degradation processes are the Weiner, Inverse-Gaussian, and Gamma processes (See Lawless and Crowder [6]) including general path model. The general path model, introduced by Lu and Meeker [2] is widely used to model degradation data. Ye and Xie [7] have done a comparative study between general path model and stochastic processes for modeling degradation data.
@@
Estimation of residual lifetime distribution of systems operating in the field plays a key role in implementing condition-based maintenance decision making (see Jardine and Banjevic [8]). Different degradation models are considered to derive the distribution function of residual lifetime. Gebraeel et al. have introduced ([9], [10]) a Bayesian degradation path model for predicting residual life time of units. Zhou et al. [11] have proposed an empirical Bayes approach to update the stochastic parameters of the degradation model for predicting soft-failure of a functioning device. Liu et al. [12] have considered modified Weiner process to predict the remaining useful lifetime; and Si et al. [13] have introduced a Wiener-process-based degradation model with a recursive filter algorithm for remaining useful lifetime prediction. A key assumption in most of the degradation models is that the degradation rate is homogeneous, see for example, Lu and Meeker [2] and Robinson and Crowder [14]. However, in practice, the assumption of homogeneity is not appropriate, particularly in presence of high variability in the manufacturing process, where the population consists of a number of homogeneous sub-populations. Ye et al. [15] have considered heterogeneous degradation rates for estimating the distribution of remaining useful life. Wen et al. [16] considered Wiener process to predict remaining lifetime in presence of heterogeneous population.
@@
Different mixture models are used for modeling heterogeneous failure time data. Mixture of gamma distributions and mixture of normal distributions (Kontar et al. [17]) are commonly used for mixture models. In the context of degradation data, Yuan and Ji [18] analyzed Laser emitters data and noticed that some units degrade faster than other units indicating the fact that they come from different sub-populations. They have considered a finite mixture of normal distributions for random effects in the linear degradation model. Model selection criteria like Akaike information criterion (AIC) and Bayesian information criterion (BIC) are used to estimate the number of mixture components. One of the major difficulty in using finite mixture models is defining the true number of components in the mixture, which denotes different sub-population under study. Limiting true number of mixture components may lead to wrong estimate of parameters. To overcome the above limitation of pre-specifying the number of sub-populations, Bayesian nonparametric models are considered in recent years.
@@
Dirichlet process introduced by Fergusan (See Fergusan [19]) is one of the most popular Bayesian nonparametric methods. Dirichlet process prior puts probability 1 to the set of discrete probability measures. To get rid of this problem Dirichlet process mixture (DPM) of normal distributions is used in the literature (See Escobar and West [20]), when the true distribution is continuous. Lo [21] introduced Monte Carlo simulation based methods for density estimation using Dirichlet process mixtures. Escobar ([22], [20]) and MacEachern [23] developed Gibbs sampling methods for Dirichlet process in a normal mixture model. But these techniques tended to produce limited posterior inference for example the Markov chain produced by these techniques tends to mix very slowly. To overcome this problem, Ishwaran and Zarepour ([25], [24]) introduced Gibbs sampling methods for the approximate Dirichlet process which consider a truncation approximation of Dirichlet process. In the context of modeling heterogeneous degradation data, Bayesian non-parametric method have gained popularity in recent times, see for example see Santos and Loschi [26], Li et al. [27] and Cheng and Yuan [28]. Nguyen et al. [29] used Bayesian non-parametric model for remaining useful life prediction of individual units with sparse degradation data.
@@
In this article, we introduce a simulation based prediction method for residual lifetime of a functioning unit, when the unit comes from a heterogeneous population. We estimate residual lifetime distribution of a functioning unit conditioned on the historical degradation measures of other units and degradation observation of the concerned unit. Firstly, we consider a Bayesian semi-parametric technique, where the degradation path is modeled using general path model and the distribution of random effect is modeled by Dirichlet process mixture of normal distribution, and parametric prior is considered for other model parameters. Utilizing the idea of Ishwaran et al. [24], we propose a hierarchical model and generate samples from posterior distributions of parameters by Gibbs sampling method. We show that the residual lifetime distribution depends on the posterior samples of parameters and it does not have closed form expression. So an approximate residual lifetime distribution is derived and finally transformation based MCMC technique [30] is used for simulating samples from the approximated distribution. Finally, we use the simulated samples for predicting residual life time of a unit. We compare the accuracy of predicted results produced by our proposed Bayesian semi-parametric method with Bayesian parametric method, where a parametric distribution is considered for the random effect and other parameters.
@@
The rest of this paper is organized as follows. Degradation model is discussed in Section 2. Residual lifetime distribution is obtained based on proposed degradation model in Section 3. We discuss method of generating observations from posterior distribution in Section 4. A simulation study is undertaken to asses the performance and efficiency of the proposed Bayesian semi-parametric method in Section 5. A real-life data on fatigue crack size is analyzed to illustrate the proposed methodology in Section 6. We conclude this article in section 7.
@@ 2.1 Bayesian semi-parametric degradation model
Let $Y_{t}$ be random variable for degradation measurement of a unit at the $t$th time point. Consider the general path model introduced by Meeker [2], where $Y_{t}$ is assumed to have the following representation.
$Y_{t}=\eta(t;\bm{\alpha},\bm{\beta})+\epsilon_{t}\hskip 11.38109pt\forall\hskip 2.27621ptt\geq 0,$ (1)
@@
where $\eta(\cdot;\bm{\alpha},\bm{\beta})$ denotes the true degradation path, $\bm{\alpha}=(\alpha_{1},\alpha_{2},.....\alpha_{k_{1}})^{t}$ is a $k_{1}$ 1 vector of parameters that are common to all units, known as fixed effect and $\bm{\beta}=(\beta_{1},\beta_{2},.....,\beta_{{k_{2}}})^{t}$ is a $k_{2}$ 1 vector of random effects for representing individual unit characteristics. Further it is assumed that the measurement errors $\epsilon_{t}$ are i.i.d $\mathcal{N}(0,\sigma_{\epsilon}^{2})$, $\forall\hskip 2.27621ptt$. It is assumed that the random effect $\bm{\beta}$ is independent of $\epsilon_{t}$, for $\forall\hskip 2.27621ptt\geq 0$. Suppose we have $n$ sample units randomly selected from the population and degradation measurements are taken at time points $t_{i1},t_{i2},...,t_{in_{i}}$ for $i$th individual where $i=1,......n$. Let $Y_{ij}$ be random variable for the degradation measurement of $i$th unit at time $t_{ij}$. Using equation $(1)$, $Y_{ij}$ is represented as follows.
[…]
@@
For each sample unit $i$, the degradation path is defined to be an observed sequence of degradation readings $\textbf{y}_{i}$ over time $\textbf{t}_{i}$, where $\textbf{y}_{i}=(y_{i1},...,y_{i{n_{i}}})$, $\textbf{t}_{i}=(t_{i1},...,t_{in_{i}})$ and $\bm{\beta}_{i}$ is the random effect for unit $i$. Linear and nonlinear structures for true degradation path in general path models are considered in different contexts. For example Lu and Meeker [2] have considered a non-linear model to estimate reliability of a unit. They considered that a unit fails when the true path $\eta(\cdot)$ crosses some predetermined threshold value, say $D$. On the other hand, Robinson and Crowder [14], Zhou et al. [11], Gebraeel et al. [10], assumed that failure of a unit occurs if its observed degradation measurement reaches a predetermined threshold $D$. This threshold value may be fixed or may vary with time.
@@
One of the important issues in using the general path models is the specification of the distribution of the random effects. The random-effects $\bm{\beta}$ represents unit wise effect and can be interpreted as degradation rates. Most of the existing literature assumes that the units under test originate from homogeneous populations, hence a unimodal distribution is assumed for the random effects for modeling degradation path. But in practice, it may happen that units come from heterogeneous population consisting of different homogeneous sub-populations. In this context, Yuan and Ji [18] have used finite mixture distribution for random effects to incorporate heterogeneity. Bayesian method is considered to account for uncertainty of number of mixture components. But in finite mixture modeling, one of the major problems is that the model assumes a finite mixture of $K$ component regardless of sample size $n$ and that ignores the fact that $K$ can also grow as the sample size $n$ increases. It may happen that a unit is degrading significantly different from all other units and a finite mixture model will not able to recognize this unit as it is coming from distinct sub-population other than $K$ many components chosen for modeling. This may lead to biased parameter estimates which in turn effects the estimation of reliability and residual lifetime distribution. To deal with this situation infinite mixture model can be considered (for example see Santos and Loschi [26]), where the number of mixture components slowly grows with sample sizes. A Bayesian alternative is to consider unknown distribution function $F$ for the corresponding random variable and hence construction of a prior for $F$ is necessary. This will also help to avoid misspecification of the distribution. In this article we consider a unknown distribution function for random effects to overcome the above-mentioned problems.
@@
Let us consider $\textbf{Y}_{train}$ = $\{Y_{ij}\}_{i=1,...,n}^{j=1,...,n_{i}}$ as the random variable that denotes the degradation measures of $n$ units. We consider these $n$ units as training set. Suppose that there is a new unit with degradation observations at some time points $t_{1},....,t_{k}$, where $t_{1}<....<t_{k}$. Let us consider the new unit as $(n+1)$th unit, where $\textbf{Y}_{n+1}=\textbf{Y}_{new}=\{Y_{new,t_{1}},....,Y_{new,t_{k}}\}$ is the random variable for degradation measurements of the new unit. Let us consider $\{y_{new,t_{1}},....,y_{new,t_{k}}\}$ as the observed degradation measures of the new unit and $\textbf{t}_{n+1}=\{t_{1},...,t_{k}\}$. We consider the random variable $\bm{Y}=[\textbf{Y}_{train},\textbf{Y}_{new}]$, which represents the degradation measures of the units in the training set and the new unit. We utilize the entire degradation information contained in the $(n+1)$ units to model the degradation path which helps in deriving the residual lifetime distribution of the new unit. In this work we use mixed effect linear model introduced by Lu and Meeker [2] for degradation modeling with one fixed effect $\alpha$ and random effect $\beta_{i}$, $i=1,...(n+1)$. Let us consider $\beta_{(n+1)}$ be the random effect on this new unit, which we represent as $\beta_{new}$. We assume that the random effect $\beta_{i}|F\overset{\mathrm{iid}}{\sim}F$, $i=1,...,(n+1)$, where $F$ is unspecified distribution and has a density function. The proposed mixed effect model is given as:
$Y_{ij}$ $=\alpha+\beta_{i}t_{ij}+\epsilon_{ij},\hskip 5.69054pti=1,....,(n+1),j=1,....,n_{i}$ (3)
[…]
@@
Next we consider selection of prior for $F$. One of the commonly used priors for $F$ is the Dirichlet process (DP) prior that is a random probability measure defined on the space of distribution functions, introduced by Fergusan [19]. Dirichlet process prior is almost surely (a.s) discrete hence it would be inappropriate to use this since we assumed that $F$ has a density function. In order to overcome this problem, we consider Dirichlet process mixture (DPM) of continuous distributions (See Escobar [20], Lo [21]). Also we consider a prior $\pi_{1}$ on the fixed effect parameter $\alpha$ and the $\pi_{2}$ on variance of measurement error $\sigma_{\epsilon}^{2}$ and they are independent to each other. In the proposed method, the degradation path is modelled using general path model and distribution of random effect is modelled through Bayesian non-parametric method, while parametric prior is considered for other model parameters. So essentially we consider a Bayesian semi-parametric technique for our analysis.
@@
In a Dirichlet process mixture modeling, the distribution function of random effect is represented by mixture over some simple parametric distribution functions where the mixing distribution is given by a Dirichlet process prior. We consider that $\beta_{i}|\theta_{i}\sim f(\cdot)$, $i=1,...,(n+1)$, where $f(\cdot)$ is a continuous distribution on $\mathbb{R}$ and $\theta_{i}$’s are unobserved random elements. Suppose that $\theta_{i}|G\overset{\mathrm{iid}}{\sim}G$ and G has a Dirichlet process prior denoted by $G\sim DP(\gamma,G_{0})$, where $\theta_{i}$in $\Theta$, $i=1,...,(n+1)$ and $G_{0}$ is the center or baseline probability measure on the measurable space ($\Theta$$\mathscr{B}$), where $\mathscr{B}$ is the corresponding Borel $\sigma$-algebra and $\gamma\in R^{+}$, considered as concentration parameter. Under these assumptions, for $i=1,.....,(n+1)$, and $j=1,....,n_{i}$, the proposed degradation model can be hierarchically represented as :
@@
${\beta}_{i}|\theta_{i}\overset{\mathrm{ind}}{\sim}f(\cdot)$

### v2

@@ Abstract
Degradation data are considered for assessing reliability in highly reliable systems. The usual assumption is that degradation units come from a homogeneous population. But in presence of high variability in the manufacturing process, this assumption is not true in general; that is different sub-populations are involved in the study. Predicting residual lifetime of a functioning unit is a major challenge in the degradation modeling, especially in heterogeneous environment. To account for heterogeneous degradation data, we have proposed a Bayesian semi-parametric approach to relax the conventional modeling assumptions. We model the degradation path using the Dirichlet process mixture of normal distributions. Based on the samples obtained from posterior distribution of model parameters we obtain residual lifetime distribution for individual unit. Transformation based MCMC technique is used for simulating values from the derived residual lifetime distribution for prediction of residual lifetime. A simulation study is undertaken to check performance of the proposed semi-parametric model compared with parametric model. Fatigue Crack Size data is analyzed to illustrate the proposed methodology.
@@ 1 Introduction
For some highly reliable engineering products, it is difficult to estimate reliability due to the fact that products take too long time to fail. In this context, degradation measures taken over time are used to estimate reliability. The evolution of the degradation measures may be observed using sensor technology through a procedure known as Condition Monitoring (See Nelson [2]). Some examples of degradation data include vibration signals for monitoring excessive wear induced in rotating machinery, acoustic emissions for monitoring crack propagation, temperature changes and oil debris for engine lubrication, decrease of brightness of light bulbs, etc. Inferences on lifetime distribution and residual lifetime of a product can be done by modeling the underlying degradation mechanism that represents the evolution of degradation resulting failure. There have been a number of works on degradation modeling, see for example, Lu and Meeker [3], Padgett and Tomlinson [4], Muller and Zhang [5] and Park and Padgett [6]. Degradation data provide more information than the bare lifetimes and hence precise inferences can be made as discussed in Lu and Meeker [3]. The modeling of degradation data can be considered by different processes. The commonly used degradation processes are the Wiener, Inverse-Gaussian, and Gamma processes (See Lawless and Crowder [7]). Lu and Meeker introduced general path model [3] to model degradation data. Ye and Xie [9] have done a comparative study between general path model and stochastic processes for modeling degradation data.
@@
Estimation of residual lifetime distribution of systems operating in the field plays a key role in implementing condition-based maintenance decision making (see Jardine and Banjevic [10]). Different degradation models are considered to derive the distribution function of residual lifetime. Gebraeel et al. ([11], [12]) have introduced a Bayesian degradation path model for predicting residual life time of units. Zhou et al. [13] have proposed an empirical Bayes approach to update the stochastic parameters of the degradation model for predicting soft-failure of a functioning device. Liu et al. [14] have considered modified Weiner process to predict the remaining useful lifetime; and Si et al. [16] have introduced a Wiener-process-based degradation model with a recursive filter algorithm for remaining useful lifetime prediction. Recently, Qin et al. [23] proposed a sparse variational Bayesian technique for remaining useful life prediction. Xu et al. [15] developed a robust Bayesian framework to estimate degradation state for remaining useful lifetime prediction in the presence of outliers. Brumm et al. [25] considered a joint model of degradation signals and time‐to‐event data for the prediction of remaining useful life. Qin and Shen [24] introduced functional variance process for degradation modeling and remaining useful lifetime prediction. In practice it is often seen that different units degrade at different rates which motivates to incorporate heterogeneity in degradation modeling ( for example see, Wang et al. [21], Zhai et al. [22]). There can be several reasons for heterogeneity, such as unobservable factors, unobservable usage patterns ( see Ye et al. [8]). A key assumption in most of the degradation models is that the degradation rate is homogeneous, see for example, Lu and Meeker [3] and Robinson and Crowder [17]. However, in practice, the assumption of homogeneity is not appropriate, particularly in presence of high variability in the manufacturing process, where the population consists of a number of homogeneous sub-populations. Ye et al. [18] have considered heterogeneous degradation rates to estimate the distribution of remaining useful life. Xu et al. [20] considered product to product heterogeneity to analyze fatigue-crack data. Wen et al. [27] considered Wiener process to predict remaining lifetime in presence of heterogeneous population. Xu et al. [28] considered a recursive Bayesian method for remaining useful life prediction for gamma degradation process with heterogeneous effects under conjugate priors.
@@
Different mixture models are used for modeling heterogeneous failure time data. Mixture of gamma distributions and mixture of normal distributions (Kontar et al. [29]) are commonly used for mixture models. In the context of degradation data, Yuan and Ji [30] analyzed Laser emitters data and noticed that some units degrade faster than other units indicating the fact that they come from different sub-populations. They have considered a finite mixture of normal distributions for random effects in the linear degradation model. Model selection criteria like Akaike information criterion (AIC) and Bayesian information criterion (BIC) are used to estimate the number of mixture components. One of the major difficulty in using finite mixture models is defining the true number of components in the mixture, which denotes different sub-population under study. Limiting true number of mixture components may lead to wrong estimate of parameters. To overcome the above limitation of pre-specifying the number of sub-populations, Santos and Loschi [39] considered Bayesian nonparametric models for degradation data.
@@
Dirichlet process introduced by Fergusan [31] is one of the most popular Bayesian nonparametric methods. Dirichlet process prior puts probability 1 to the set of discrete probability measures. To get rid of this problem Dirichlet process mixture (DPM) of normal distributions is used in the literature (See Escobar and West [32]), when the true distribution is continuous. Lo [33] introduced Monte Carlo simulation based methods for density estimation using Dirichlet process mixtures. Escobar ([34], [32]) and MacEachern [35] developed Gibbs sampling methods for Dirichlet process in a normal mixture model. But these techniques tended to produce limited posterior inference for example the Markov chain produced by these techniques tends to mix very slowly. To overcome this problem, Ishwaran and Zarepour ([37], [36]) introduced Gibbs sampling methods for the approximate Dirichlet process which consider a truncation approximation of Dirichlet process. In the context of modeling heterogeneous degradation data, Bayesian non-parametric method have gained popularity in recent times, see for example, Santos and Loschi [39], Li et al. [41] and Cheng and Yuan [42]. Recently Li et al. [26] used non-parametric techniques to model heterogeneous degradation data. Nguyen et al. [43] used Bayesian non-parametric model for remaining useful life prediction of individual units with sparse degradation data, but they consider Dirichlet process prior for distribution of model parameters which neglects the fact that the true distribution of the parameters for random effect in the hierarchical model may be continuous. Bayesian semi-parametric techniques are widely recognized for mixed effect models due to its flexibility to model the multi-modal, skewed nature of random effect distribution ( see Ishwaran and Takahara [44], Zhang and Davidian [45]). There is lack of research on modeling degradation data using Bayesian semi-parametric techniques to study the distributions of residual lifetimes of individual units. Also, there are not much literature on comparison of Bayesian parametric and semi-parametric method in the context of degradation modeling, which is essential to understand the validity of model assumptions. Furthermore, the distribution of residual lifetime is required to properly characterized, which is essential for prediction.
@@
In this article, we introduce a simulation based prediction method for residual lifetime of a functioning unit, when the unit comes from a heterogeneous population. The main contributions of this article are summarized as follows. We consider a Bayesian semi-parametric technique, where the degradation path is modeled using general path model and the distribution of random effect is modeled by Dirichlet process mixture of normal distribution, and parametric prior is considered for other model parameters. The proposed semi-parametric construction deals with the problem when the true random effect distribution is continuous. We estimate residual lifetime distribution of a functioning unit conditioned on the observed degradation measures of other units and degradation observation of the concerned unit. Samples generated from posterior distributions of parameters by Gibbs sampling method. We show that the residual lifetime distribution depends on the posterior samples of parameters and it does not have closed form expression. So an approximate residual lifetime distribution is derived. We use transformation based MCMC technique introduced by Dutta and Bhattacharya [46] to simulate samples from the approximated distribution. Finally, we use the simulated samples for predicting residual life time of a unit. The accuracy of predicted results produced by the proposed Bayesian semi-parametric method is compared with Bayesian parametric method, where a unimodal parametric distribution is considered for the random effect and other parameters.
@@
The rest of this paper is organized as follows. Degradation model is discussed in Section 2. Residual lifetime distribution is obtained based on proposed degradation model in Section 3. We discuss method of generating observations from posterior distribution in Section 4. Prediction evaluation criteria to check the performance of the proposed models are given in Section 5. A simulation study is undertaken to asses the performance and efficiency of the proposed Bayesian semi-parametric method in Section 6. A real-life data on fatigue crack size is analyzed to illustrate the proposed methodology in Section 7. We conclude this article in Section 7.
@@ 2.1 Bayesian semi-parametric degradation model
Let $Y_{t}$ be the random variable denoting the degradation measurement of a unit at the $t$th time point. Consider the general path model introduced by Lu and Meeker [3], where $Y_{t}$ is assumed to have the following representation.
$Y_{t}=\eta(t;\bm{\alpha},\bm{\beta})+\epsilon\hskip 11.38109pt$ (1)
@@
where $\eta(\cdot;\bm{\alpha},\bm{\beta})$ denotes the true degradation path, $\bm{\alpha}=(\alpha_{1},\alpha_{2},\dots,\alpha_{k_{1}})^{t}$ is a $k_{1}$ × 1 vector of parameters that are common to all units, known as fixed effect and $\bm{\beta}=(\beta_{1},\beta_{2},\dots,\beta_{{k_{2}}})^{t}$ is a $k_{2}$ × 1 vector of random effects for representing individual unit characteristics. Further it is assumed that the measurement error at the $t$th time point is $\epsilon$ and $\epsilon\sim\mathcal{N}(0,\sigma_{\epsilon}^{2})$. It is assumed that the measurement errors for distinct time points are identically distributed and independent to each other. We also assume that random effect $\bm{\beta}$ is independent of $\epsilon$. Suppose we have $n$ sample units randomly selected from the population and degradation measurements are taken at time points $t_{i1},t_{i2},\dots,t_{in_{i}}$ for $i$th individual where $i=1,\dots,n$. Let $Y_{ij}$ be random variable denoting the degradation measurement of $i$th unit at time $t_{ij}$. Using equation $(1)$, $Y_{ij}$ is represented as follows.
$Y_{ij}=\eta(t_{ij};\bm{\alpha},\bm{\beta}_{i})+\epsilon_{ij},\hskip 11.38109pti=1,\dots,n,j=1,\dots,n_{i},$ (2)
@@
For each sample unit $i$, the degradation path is defined to be an observed sequence of degradation readings $\textbf{y}_{i}$ over time $\textbf{t}_{i}$, where $\textbf{y}_{i}=(y_{i1},\dots,y_{i{n_{i}}})$, $\textbf{t}_{i}=(t_{i1},\dots,t_{in_{i}})$ and $\bm{\beta}_{i}$ is the random effect for unit $i$. Linear and nonlinear structures for true degradation path in general path models are considered in different contexts. For example Lu and Meeker [3] have considered a non-linear model to estimate reliability of a unit. They considered that a unit fails when the true path $\eta(\cdot)$ crosses some predetermined threshold value, say $D$. On the other hand, Robinson and Crowder [17], Zhou et al. [13], Gebraeel et al. [12], assumed that failure of a unit occurs if its observed degradation measurement reaches a predetermined threshold $D$. This threshold value may be fixed or may vary with time.
@@
One of the important issues in using the general path models is the specification of the distribution of the random effects. The random-effects $\bm{\beta}$ represent unit wise effect and can be interpreted as degradation rates. Most of the existing literature assumes that the units under test originate from homogeneous populations, hence a unimodal distribution is assumed for the random effects for modeling degradation path. But in practice, it may happen that units come from heterogeneous population consisting of different homogeneous sub-populations. In this context, Yuan and Ji [30] have used finite mixture distribution for random effects to incorporate heterogeneity. Bayesian method is considered to account for uncertainty of number of mixture components. But in finite mixture modeling, one of the major problems is that the model assumes a finite mixture of $K$ component regardless of sample size $n$ and that ignores the fact that $K$ can also grow as the sample size $n$ increases. It may happen that a unit is degrading significantly different from all other units and a mixture model with fixed number of mixing components will not able to recognize this unit as it is coming from distinct sub-population other than $K$ many components chosen for modeling. This may lead to biased parameter estimates which in turn effects the estimation of reliability and residual lifetime distribution. A flexible approach to deal with this problem is to consider a unknown distribution function for random effects (see Dunson [38]). A Bayesian approach to this problem is done by considering unknown distribution function $F$ as a random quantity, and hence construction of a prior for $F$ is necessary. To deal with this situation Bayesian nonparametric techniques can be considered ( see Santos and Loschi [39]), where the number of mixture components slowly grows with sample sizes. This will also help to avoid misspecification of the distribution. In this article, we consider an unknown distribution function for random effects to overcome the above-mentioned problems.
@@
Let us consider $\textbf{Y}_{train}$ = $\{Y_{ij}\}_{i=1,\dots,n}^{j=1,\dots,n_{i}}$ as the collection of random variables that denotes the degradation measures of $n$ units. We consider these $n$ units as training set. Suppose that there is a new unit with degradation observations at some time points $t_{1},\dots,t_{k}$, where $t_{1}<\dots<t_{k}$. Let us consider the new unit as $(n+1)$th unit, where $\textbf{Y}_{n+1}=\textbf{Y}_{new}=\{Y_{new,t_{1}},\dots,Y_{new,t_{k}}\}$ is the collection of random variables for degradation measurements of the new unit. Let us consider $\textbf{y}_{new}=\{y_{new,t_{1}},\dots,y_{new,t_{k}}\}$ as the observed degradation measures of the new unit and $\textbf{t}_{n+1}=\{t_{1},\dots,t_{k}\}$. We consider $\bm{Y}=[\textbf{Y}_{train},\textbf{Y}_{new}]$, which is the collection of random variables that represents the degradation measures of the units in the training set and the new unit. We utilize the entire degradation information contained in the $(n+1)$ units to model the degradation path which helps in deriving the residual lifetime distribution of the new unit. In this work we use mixed effect model introduced by Lu and Meeker [3] for degradation modeling with one fixed effect $\alpha$ and random effect $\beta_{i}$, $i=1,\dots,(n+1)$. Let us consider $\beta_{(n+1)}$ be the random effect on this new unit, which we represent as $\beta_{new}$. We assume that the random effect $\beta_{i}|F\overset{\mathrm{iid}}{\sim}F$, $i=1,\dots,(n+1)$, where $F$ is unspecified distribution and has a density function. The proposed mixed effect model is given as:
$Y_{ij}$ $=\eta(t_{ij};\bm{\alpha},\bm{\beta}_{i})+\epsilon_{ij},\hskip 5.69054pti=1,\dots,(n+1),j=1,\dots,n_{i}$ (3)
$\beta_{i}|F\overset{\mathrm{iid}}{\sim}F,\hskip 5.69054pt\epsilon_{ij}\overset{\mathrm{iid}}{\sim}\mathcal{N}(0,\sigma_{\epsilon}^{2}),\hskip 5.69054pt\epsilon_{ij}\perp\beta_{i}\hskip 5.69054pt\text{for all $i,j$}$
@@
It is important to consider a proper structure for $\eta(t_{ij};\bm{\alpha},\bm{\beta}_{i})$, given in equation (3). Considering a linear degradation model can often help in deriving closed form expression for posterior distribution, but one can also consider non-linear models instead of linear model. The traditional approach is to compare those models using model selection criteria like Bayesian Information Criterion (BIC), Conditional Predictive Ordinate (CPO), Deviance Information Criterion (DIC) etc. and choose the best model. In this article, we conduct analysis based on the linear degradation model where, $\eta(t;\bm{\alpha},\bm{\beta})=\alpha+\beta t$ $\forall\hskip 2.84526ptt>0$. Next we consider selection of prior for $F$. One of the commonly used priors for $F$ is the Dirichlet process (DP) prior, which is a probability measure defined on the space of distribution functions, introduced by Fergusan [31]. Dirichlet process prior is almost surely (a.s) discrete hence it would be inappropriate to use this since we assumed that density function exits with respect to $F$. In order to overcome this problem, we consider Dirichlet process mixture (DPM) of continuous distributions (See Escobar [32], Lo [33]). Also we consider a parametric prior $\pi_{1}(\cdot)$ on the fixed effect parameter $\alpha$ and the $\pi_{2}(\cdot)$ on variance of measurement error $\sigma_{\epsilon}^{2}$ and they are independent to each other. In the proposed method, the degradation path is modeled using general path model and distribution of random effect is modeled through Bayesian non-parametric method, while parametric prior is considered for other model parameters. So essentially we consider a Bayesian semi-parametric technique for our analysis.
@@
In a Dirichlet process mixture modeling, the distribution function of random effect is represented by mixture over some simple parametric distribution functions where the mixing distribution is given by a Dirichlet process prior. We consider that $\beta_{i}|\theta_{i}\sim f(\cdot)$, $i=1,\dots,(n+1)$, where $f(\cdot)$ is a continuous distribution on $\mathbb{R}$ with a probability density function and $\theta_{i}$’s are unobserved random elements. Suppose that $\theta_{i}|G\overset{\mathrm{iid}}{\sim}G$ and G has a Dirichlet process prior denoted by $G\sim DP(\gamma,G_{0})$, where $\theta_{i}$ $\in$ $\Theta$, $i=1,\dots,(n+1)$ and $G_{0}$ is the center or baseline probability measure on the measurable space ($\Theta$$\mathscr{B}$), where $\mathscr{B}$ is the corresponding Borel $\sigma$-algebra and $\gamma\in\mathbb{R}^{+}$, considered as concentration parameter. Under these assumptions, for $i=1,\dots,(n+1)$, and $j=1,\dots,n_{i}$, the proposed degradation model can be hierarchically represented as :
@@
${\beta}_{i}|\theta_{i}\overset{\mathrm{ind}}{\sim}f(\beta_{i}|\theta_{i})$